differential equations

Problem 3 Solve V + 3y + 2y = uz(t).
Answer #1

Angi differential equation We have to solve Y+ 3y + 2y = U2 (A) for (if We take y+3yltay zo » (D?+ 3D+2) y =0 m² +3m +2=0 m